OpenStudy (anonymous):

Can someone please help me ? Im trying to factor this equation. F(x)= (-x)^2 - 10x -3 Dont just give me the answer, i need explanation please. I'm doing a COMPASS test on weds

OpenStudy (anonymous):

\[-x^{2}-10x-3 \rightarrow x=5\pm \sqrt{25-3}\] \[f(x)= (x-5-\sqrt{22})(x-5+\sqrt{22})\]

OpenStudy (anonymous):

how do you get the square root of 22 ? Can you make it an easier explanation for me because i first set it up as (x )(x ) and i couldnt figure out the rest because no factors of -3 were equal to the sum -10

OpenStudy (usukidoll):

is your equation \[-x^2-10x-3 \] ?

OpenStudy (anonymous):

there exists the formula: \[\Delta= \frac{ -\frac{ b }{ 2 }\pm \sqrt{(\frac{ b }{ 2 }})^{2}-ac }{ a }\] where a,b and c are the numbers in front of x^2, x, and the last term, respectively.

OpenStudy (mrnood):

there is a difference between \[-x ^{2} and (-x)^{2}\] which do you mean?

OpenStudy (usukidoll):

Maybe let's assume that it is -x^2 ? since the original poster did say that they couldn't factor ?!

OpenStudy (mrnood):

y= -x^2-10x-3 has 2 real roots, but is not factorisable

OpenStudy (usukidoll):

true @MrNood because when we used the discriminant formula \[b^2-4ac \] letting a = -1, b = -10, and c = -3 we don't have a perfect square

OpenStudy (mrnood):

same is true for y= (-x)^2 - 10x -3

OpenStudy (mrnood):

of course we can use the quadratic formula ot get the solution BUT the OP asks for factorisation @francemaz I think your version of the equation is not correct

OpenStudy (usukidoll):

We have no choice but to use the quadratic formula to get the roots so we can do factorization. Otherwise, what's the alternative? @MrNood

OpenStudy (usukidoll):

only part of his version is correct. There are sign errors

OpenStudy (mrnood):

using the quadratic formula will get the 2 solutions. This is not 'factorisation' in my understanding of the term

OpenStudy (usukidoll):

then just leave the problem alone afterwards? are you serious? By the discriminant formula, we either have a perfect square (yes we can factor) or not a perfect square (we have to use the quadratic equation)

OpenStudy (anonymous):

given eq can be written as (-x)^2+ 10(-x) -3 now use quadratic formula using -x instead of x and a=1, b= 10 , c=-3,